# Tuning electronic and optical properties of 2D polymeric C60 by stacking two layers

### v1

@@ 1 Introduction
Recently, monolayer fullerene networks have been synthesised through organic cation slicing exfoliation [1]. This combines the characteristics of 0D fullerene molecules and 2D graphene monolayers such as large surface areas and high flexibility, making it promising for applications such as photocatalytic water splitting [2, 3, 4]. However, achieving stable monolayer fullerenes remains challenging due to their inherent conflicting dynamic, thermodynamic and mechanical stability [5]. This results in C60 monolayers being prone to degradation into 1D or 0D at high temperatures. As a result, it has been found that mechanical exfoliation of van der Waals polymeric C60 can only lead to few-layer covalent network of fullerenes instead of the single-layer form [6]. The obtained few-layer fullerene networks, consisting of two or more stacked layers, have been more commonly produced and studied experimentally [6, 7]. Few-layer C60 networks benefit from enhanced structural stability due to stronger interlayer van der Waals interactions. This provides more stable configurations in multiple structural phases [1, 6] while still retaining many of the desirable chemical properties predicted in monolayers such as photocatalysis [7].
@@ 2 Methods
Density functional theory (DFT) calculations were performed using the Vienna ab initio Simulation Package (VASP) [11, 12]. The projector-augmented wave (PAW) basis set [13, 14] was employed with the Perdew-Burke-Ernzerhof functional revised for solids (PBEsol) [15] within the generalised gradient approximation (GGA). A plane-wave cutoff energy of 800 eV was used with the Brillouin zone sampled using a $\Gamma$-centred k-point grid of $2\times 3$. The zero damping DFT-D3 method of Grimme [16] was included to described the van der Waals interactions. An interlayer spacing larger than 20 Å was applied with the dipole corrections [17] along $z$. The crystal structures were fully relaxed using the energy and force convergence criteria of $10^{-6}$ eV and $10^{-2}$ eV/Å respectively. For electronic structures, hybrid functional calculations was used by mixing 75% of the PBEsol exchange functional with 25% unscreened exact Hartree-Fock exchange [18], denoted as PBEsol0. The effective masses at the band edges were calculated by sampling the Brillouin Zone near $\Gamma$ in intervals of 0.015 Å-1 with 5 points each along $\Gamma-\mathrm{X}$ and $\Gamma-\mathrm{Y}$, and then fitting the energy dispersion curves with a cubic polynomial. Excitonic calculations were performed using the time-dependent Hartree-Fock (TDHF) method with the Casida equation [19] from a basis of the 16 highest valance bands and the 16 lowest conduction bands. A $\mathbf{k}$-mesh of $4\times 6$ was then used for the monolayer and $2\times 3$ for the bilayer, both of which were converged well.
@@ 3.1 Crystal structures
Figure 1 shows the crystal structures of monolayer and bilayer fullerene networks. For the C60 monolayers, each carbon cage is connected by the C--C single bonds along the [110] direction and the in-plane [$2+2$] cycloaddition bonds along the [010] direction, leading to a quasi-hexagonal lattice with space group $Pmna$ (No. 53). Bilayer fullerene networks have a space group $P2/c$ (No. 13) and can be viewed as an AB stacking of two layers, which is held together by the van der Waals interactions. The lattice constants of the bilayer $a=15.804$ and $b=9.115$ are slightly larger than those of the monolayer $a=15.798$ and $b=9.110$, agreeing well with the synthesised structures [1, 6].
@@ 3.2 Band structures
The PBEsol0 band structures of monolayer and bilayer fullerene networks are shown in Fig. 2. Both structures show a direct band gap at $\Gamma$ with the bilayer gap of 2.05 eV slightly lower than the monolayer gap of 2.08 eV, which is due to interlayer interactions in the bilayer. The computed band gap for the bilayer agrees well with the measured gap for few-layer polymeric C60 of 2.02 [6] and 2.05 eV [7], while the monolayer gap from unscreened hybrid functional PBEsol0 is comparable with the gap from the many-body perturbation theory [20]. Because the band gap of bilayer polymeric C60 is similar to that of the monolayer, the energy levels of their conduction band minimum (CBM) and valence band maximum (VBM) straddle the redox potentials of water in Fig. 3(a), similar to monolayer fullerene networks [2]. As a result, the band edges of bilayer C60 networks provide sufficient driving force for overall water splitting, as observed experimentally [7].
In bilayer fullerene networks, there are two nearly degenerate VBM states and two nearly degenerate CBM states, denoted as VBM-1/VBM and CBM/CBM+1 respectively (labelling from low to high eigenenergies). The nearly degenerate VBM-1 and VBM (CBM and CBM+1) states can be understood as the in-phase and out-of-phase combinations of the VBM (CBM) states of the top and bottom layers that make up the bilayer, as demonstrated in Fig. 3(b). This suggests similar orbital characters on the top and bottom layers, preserving the band-edge positions and hence the catalytic reactivity of the monolayers. The VBM-1 and VBM (or CBM and CBM+1) states have slightly different eigenenergies because of the interlayer interactions [21].

### v2

@@ 1 Introduction
Recently, monolayer fullerene networks have been synthesised through organic cation slicing exfoliation [1]. This combines the characteristics of 0D fullerene molecules and 2D graphene monolayers such as large surface areas and high flexibility, making it promising for applications such as photocatalytic water splitting [2, 3, 4, 5]. However, achieving stable monolayer fullerenes remains challenging due to their inherent conflicting dynamic, thermodynamic and mechanical stability [6, 7]. This results in C60 monolayers being prone to degradation into 1D or 0D at high temperatures. As a result, it has been found that mechanical exfoliation of van der Waals polymeric C60 can only lead to few-layer covalent network of fullerenes instead of the single-layer form [8]. The obtained few-layer fullerene networks, consisting of two or more stacked layers, have been more commonly produced and studied experimentally [8, 9]. Few-layer C60 networks benefit from enhanced structural stability due to stronger interlayer van der Waals interactions. This provides more stable configurations in multiple structural phases [1, 8] while still retaining many of the desirable chemical properties predicted in monolayers such as photocatalysis [9].
@@ 2 Methods
Density functional theory (DFT) calculations were performed using the Vienna ab initio Simulation Package (VASP) [13, 14]. The projector-augmented wave basis set [15, 16] was employed with the Perdew-Burke-Ernzerhof functional revised for solids (PBEsol) [17] within the generalised gradient approximation (GGA). A plane-wave cutoff energy of 800 eV was used with the Brillouin zone sampled using a $\Gamma$-centred k-point grid of $2\times 3$. The zero damping DFT-D3 method of Grimme was included to described the van der Waals interactions [18]. A vacuum spacing larger than 24 Å was applied with the dipole corrections along $z$[19]. The crystal structures were fully relaxed using the energy and force convergence criteria of $10^{-6}$ eV and $10^{-2}$ eV/Å respectively. For electronic structures, hybrid functional calculations was used by mixing 75% of the PBEsol exchange functional with 25% unscreened exact Hartree-Fock exchange [20], denoted as PBEsol0. The effective masses at the band edges were calculated by sampling the Brillouin Zone near $\Gamma$ in intervals of 0.015 Å-1 with 5 points each along $\Gamma-\mathrm{X}$ and $\Gamma-\mathrm{Y}$, and then fitting the energy dispersion curves with a cubic polynomial. Excitonic calculations were performed using the time-dependent Hartree-Fock (TDHF) method with the Casida equation [21] from a basis of the 16 highest valance bands and the 16 lowest conduction bands with a well converged $\mathbf{k}$-mesh of $4\times 6$ for the monolayer and $2\times 3$ for the bilayer. The TDHF approach is similar to the Bethe-Salpeter equation (BSE) but crucially uses the exchange-correlation kernel for the screening of the Coulomb potential instead of the screened exchange $W$, yielding quantitatively agreeable excitonic properties [2, 3] with the computationally heavy $GW$ + BSE results [22].
@@ 3.1 Crystal structures
Figure 1 shows the crystal structures of monolayer and bilayer C60 networks. For the monolayer, each carbon cage is connected by the C--C single bonds along the [110] direction and the in-plane [$2+2$] cycloaddition bonds along the [010] direction, leading to a quasi-hexagonal lattice with space group $Pmna$ (No. 53). Bilayer C60 networks have a space group $P2/c$ (No. 13) and can be viewed as an AB stacking of two layers held together by the van der Waals interactions. The AB stacking pattern leads to a closely-packed lattice as expected for nearly-spherical C60 molecules to be arranged in a space-efficient manner, as reported experimentally [8]. The lattice constants of the bilayer $a=15.804$ and $b=9.115$ are slightly larger than those of the monolayer $a=15.798$ and $b=9.110$, agreeing well with the synthesised structures [1, 8].
Table 1: Lattice constants of bilayer fullerene networks from different functionals, as well as the layer distance $\Delta h$ (defined as the height difference between the lowest atom in top layer and the highest atom in bottom layer) and the actual distance $\Delta d$ between the corresponding atoms.
$a$ (Å) $b$ (Å) $\Delta h$ (Å) $\Delta d$ (Å)
PBE 15.931 9.171 2.251 5.962
PBE + D3 15.889 9.157 1.662 5.763
PBEsol 15.840 9.128 1.912 5.885
PBEsol + D3 15.804 9.115 1.669 5.840
@@ 3.2 Band structures
The PBEsol0 band structures of monolayer and bilayer fullerene networks are shown in Fig. 2. Both structures show a direct band gap at $\Gamma$ with the bilayer gap of 2.05 eV slightly lower than the monolayer gap of 2.08 eV, which is due to the weak interlayer interactions in the bilayer. As shown in Table 1, despite that the layer distance $\Delta h$ between the lowest atom in top layer and the highest atom in bottom layer is quite small ($1.662-2.251$ for different functionals), the actual distance between the corresponding atoms $\Delta d$ is much larger (¿ 5.840 Å) because of the AB stacking, leading to the weak interlayer interactions.
The computed band gap for the bilayer agrees well with the measured gap for few-layer polymeric C60 of 2.02 [8] and 2.05 eV [9], while the monolayer gap from unscreened hybrid functional PBEsol0 is comparable with the gap from the many-body perturbation theory [22]. Because the band gap of bilayer polymeric C60 is similar to that of the monolayer, the energy levels of their conduction band minimum (CBM) and valence band maximum (VBM) straddle the redox potentials of water in Fig. 3(a), similar to monolayer fullerene networks [2]. As a result, the band edges of bilayer C60 networks provide sufficient driving force for overall water splitting, as observed experimentally [9]. In bilayer fullerene networks, there are two nearly degenerate VBM states and two nearly degenerate CBM states, denoted as VBM-1/VBM and CBM/CBM+1 respectively (labelling from low to high eigenenergies). The nearly degenerate VBM-1 and VBM (CBM and CBM+1) states can be understood as the in-phase and out-of-phase combinations of the VBM (CBM) states of the top and bottom layers that make up the bilayer, as demonstrated in Fig. 3(b). This suggests similar orbital characters on the top and bottom layers, preserving the band-edge positions and hence the catalytic reactivity of the monolayers. The VBM-1 and VBM (or CBM and CBM+1) states have slightly different eigenenergies because of the interlayer interactions [23].